OpenStudy (pleasehelpme101):

Subtract 9 y minus 39y−3 from y minus 9y−9

OpenStudy (pleasehelpme101):

so ive been studying this type of questions for a day but no matter how hard i try i can solve it :(

OpenStudy (advert):

39-9= 30 then 30 - 3 = 27 then 9-9= 0 so stays at 27 there you go easy peasy

OpenStudy (advert):

so 27 is your answer

OpenStudy (pleasehelpme101):

wow thank you so much!

OpenStudy (advert):

no problem ;)

OpenStudy (pleasehelpme101):

it was actually wrong :(